Write the slope-intercept equation for the line with slope 3/4 and y-intercept (0, -5) and graph the line
Answer by mananth(16946)

Plug value of the slope and point (0,-5) in
Y = m x + b
-5.00 = 0 + b
b= -5 0
b= -5
So the equation will be
Y = 0.75x -5
Xintercept
plug Y=0
y=mx+b
0=3/4 x -5
x=20/3
(20/3,0)


m.ananth@hotmail.ca